# Enhancing inflationary model predictions via refined slow-roll dynamics

### v2

@@ Abstract
The inflationary paradigm not only addresses early Universe puzzles but also aligns well with the observational constraints, with slow-roll inflationary models fitting best. Evaluating these model predictions requires considering both slow-roll inflationary dynamics and the subsequent reheating epoch. This involves the quantitative analysis that takes into account the effective equation of state (EoS) and duration of reheating, connecting these with the perturbations generated deep during the inflationary era. Given the complexities involved, many approximations are often used for simplification. However, as future observations are expected to improve the accuracy of these observables significantly, this work takes a different approach. Instead of relying on approximations, we consider the near-accurate solutions of inflationary models and compare our results with the pre-existing ones. It mainly incorporates two improvements: the first is the accurate dynamics of the slow-roll evolution, and, thus, the end of inflation; and the second is the higher-order slow-roll corrections to the perturbed observables. Our findings indicate that, by implementing these corrections, the theoretical predictions improve significantly. It also indicates that seemingly minor corrections can have significant effects on the perturbed observables and these refined predictions can be compared with future observations to potentially rule out models and help resolve the degeneracy problem of the inflationary paradigm.
1 Introduction
@@
Theoretical constraints derived from an inflationary model, however, not only depend on the dynamics of inflation itself but also depend on the reheating era, and therefore, studying reheating dynamics is crucial for accurately predicting the early Universe behavior. There are two main approaches to this: the first is the qualitative analysis [42, 36, 37, 38, 39, 41, 40, 43, 46] which examines the exact decay process of the inflaton field through the parametric resonance, leading to the understanding of the microphysics involved during the reheating epoch. The second approach, which is the main focus of this manuscript, is the quantitative analysis [59, 54, 60, 44, 61, 45, 47, 49, 52, 62, 63]. This approach looks at the macrophysics of the reheating epoch by constraining the duration ($N_{\rm re}$) and the equation of state (EoS) parameter ($w_{\rm re}$) during the reheating era using perturbations generated deep in the early (inflation) Universe.
@@
Improving the analysis can again be performed in two main ways. The first one involves a detailed and careful consideration of the reheating analysis that can accurately depict the reheating parameters ($N_{\rm re},w_{\rm re}$). However, since it entails dealing with various distinct microphysical processes, i.e., the qualitative analysis, we have opted not to pursue this approach as it is beyond the scope of this work. Instead, we focus on the second, more straightforward improvement: the slow-roll methodology and approximations. Recent findings in Refs. [71, 72] indicate that a near-accurate analytical solution can significantly enhance background predictions near the end of inflation, thereby improving $N_{k}$ [71]. For example, in the case of Starobinsky inflation, the improvement $\Delta N_{k}\sim 1.5$ suggests that
@@
The article is structured as follows: the following section 2 presents the model and the subsequent dynamics for the early Universe dynamics. In this section, we also demonstrate the slow-roll conditions as well as the slow-roll approximations. In the subsequent section 3, we study the quantitative analysis of reheating linking the inflationary perturbed variables to the observable constraints. This section, as an example, also examines the popular Starobinsky model of inflation, and with the help of the reheating analysis along with leading-order slow-roll approximations, we carefully obtain the observational consequences of the model. In Sec. 4, we highlight the dire need for the corrections to the dynamics, subsequently implement these changes, and obtain the improved constraint of the model, which turns out to be significant. Finally, we conclude our work in Sec. 5.
A brief explanation about some terms and symbols we’ll be using. In our work, we use natural units where $\hbar=c=k_{B}=1$. We also define the reduced Planck mass as $M_{\rm pl}\equiv(8\pi G)^{-1/2}=1$. The metric signature we use is (,+,+,+)(-,+,+,+)( - , + , + , + ). Greek indices are contracted with the metric tensor $g_{\mu\nu}$, while Latin indices will be contracted with the Kronecker delta $\delta_{ij}$. We use symbols $\partial$ and $\nabla$ for partial and covariant derivatives, respectively. Additionally, the dots and primes above symbols represent derivatives for cosmic and conformal time, respectively, related to the Friedmann-Lemaître-Robertson-Walker (FLRW) line-element.
2 General equations and slow-roll conditions
@@
$S=\frac{1}{2}\int{\rm d}^{4}x\sqrt{-g}(R-g^{\mu\nu}\partial_{\mu}\phi\partial_{\nu}\phi-2V(\phi)).$ (2.1)
@@
$T_{\mu\nu{(\phi)}}=\partial_{\mu}\phi\ \partial_{\nu}\phi-g_{\mu\nu}\left(\frac{1}{2}\partial_{\lambda}\phi\ \partial^{\lambda}\phi+V(\phi)\right).$ (2.4)
@@
${\rm d}s^{2}=-{\rm d}t^{2}+a^{2}(t){\rm d}\mathbf{x}^{2},$ (2.5)
where, $a(t)$ is the scalar factor, Eqs. (2.2) and (2.3) can be reduced to
$3H^{2}=\frac{1}{2}\dot{\phi}^{2}+V(\phi),$ (2.6)
$\dot{H}=-\frac{1}{2}\dot{\phi}^{2},$ (2.7)
$\ddot{\phi}+3H\dot{\phi}+V_{,\phi}=0.$ (2.8)
@@
$H=\sqrt{\frac{V}{3-\frac{1}{2}\phi_{N}^{2}}},\qquad\frac{H_{N}}{H}=-\frac{1}{2}\phi_{N}^{2},$ (2.9)
$\phi_{NN}+\left(3-\frac{1}{2}\phi^{2}_{N}\right)\left(\phi_{N}+\frac{V_{\phi}}{V}\right)=0,$ (2.10)
@@
$\mathcal{P}_{\mathcal{R}}=\frac{k^{3}}{2\pi^{2}}|\mathcal{R}_{k}|^{2}\equiv A_{\mathcal{R}}\left(\frac{k}{k_{*}}\right)^{n_{s}-1},\quad\mathcal{P}_{T}=2.\frac{k^{3}}{2\pi^{2}}|h_{k}|^{2}\equiv A_{T}\left(\frac{k}{k_{*}}\right)^{n_{T}}$ (2.11)
@@
$\epsilon_{\rm{1}}\equiv-\frac{\dot{H}}{H^{2}}=\frac{1}{2}\phi_{N}^{2},\qquad\epsilon_{\rm{2}}\equiv\frac{\dot{\epsilon_{\rm{1}}}}{H\epsilon_{\rm{1}}}=2\frac{\phi_{NN}}{\phi_{N}}.$ (2.12)
@@
$\epsilon_{\rm{1}}\ll 1,\quad\epsilon_{\rm{2}}\ll 1.$ (2.13)
These two condition are also referred to as slow-roll conditions. Under this conditions, Eqs. (2.9) and (2.10) become
$H^{2}\simeq\frac{1}{3}V,\quad\phi_{N}\simeq-\frac{V_{\phi}}{V},$ (2.14)
@@
$\epsilon_{1}\simeq\frac{1}{2}\left(\frac{V_{\phi}}{V}\right)^{2},\quad\epsilon_{2}\simeq 2\left(\frac{V^{2}_{\phi}}{V^{2}}-\frac{V_{\phi\phi}}{V}\right).$ (2.15)
Eqs. (2.14) are referred to as the slow-roll equations. By solving these simplified equations, one can obtain the approximate analytical solution of the Hubble parameter as well as the slow-roll parameters, and thus the dynamics of the Universe in slow-roll regime.
@@
$A_{\mathcal{R}}\simeq\frac{H^{2}}{8\pi^{2}\epsilon_{1}},~{}~{}~{}~{}\quad n_{s}\simeq 1-2\epsilon_{1}-\epsilon_{2},~{}~{}~{}~{}\quad r\simeq 16\epsilon_{1}.$ (2.16)
@@
3 Quantitative analysis of reheating
@@
To quantify the effect of reheating, let us first simplify it by parametrizing the reheating epoch using an effective EoS parameter $w_{\rm re}$, such that the energy density during reheating effectively behaves as
$\rho\propto a^{-3(1+w_{\rm re})}.$ (3.1)
@@
$N_{\rm re}\equiv{\rm ln}\left(\frac{a_{\rm re}}{a_{\rm end}}\right)$ (3.2)
@@
$\frac{k}{a_{0}H_{0}}=\frac{a_{k}H_{k}}{a_{0}H_{0}}=\frac{a_{k}}{a_{\rm end}}\frac{a_{\rm end}}{a_{\rm{re}}}\frac{a_{\rm{re}}}{a_{\rm{eq}}}\frac{a_{\rm{eq}}}{a_{0}}\frac{H_{k}}{H_{\rm{eq}}}\frac{H_{\rm{eq}}}{H_{0}},$ (3.3)
@@
${\rm ln}\left(\frac{k}{a_{0}H_{0}}\right)=-N_{k}-N_{\rm re}-N_{\rm{RD}}+{\rm ln}\left(\frac{a_{\rm{eq}}H_{\rm{eq}}}{a_{0}H_{0}}\right)+{\rm ln}\left(\frac{H_{k}}{H_{\rm{eq}}}\right),$ (3.4)
@@
$g_{\rm{s,re}}\,T_{\rm{re}}^{3}=\left(\frac{a_{0}}{a_{\rm{re}}}\right)^{3}\left(2T_{0}^{3}+6\times\frac{7}{8}T_{\nu 0}^{3}\right),$ (3.6)
@@
$\frac{T_{\rm re}}{T_{0}}=\frac{a_{0}}{a_{\rm{eq}}}\frac{a_{\rm{eq}}}{a_{\rm{re}}}\left(\frac{43}{11g_{\rm{s,re}}}\right)^{1/3},$ (3.7)
and the Eq. (3.4) can be re-written as
@@
$T_{\rm{re}}=\frac{a_{0}T_{0}}{k}\left(\frac{43}{11g_{\rm{s,re}}}\right)^{1/3}H_{k}e^{-N_{\rm k}-N_{\rm{re}}}.$ (3.9)
@@
$H_{k}=\frac{\pi\sqrt{rA_{\mathcal{R}}}}{\sqrt{2}}.$ (3.10)
@@ 3.1 Starobinsky inflation
$V(\phi)=\frac{3}{4}M^{2}\left(1-e^{-\sqrt{2/3}\phi}\right)^{2},$ (3.11)
@@
Figure 1: The Starobinsky potential given in Eq. (3.11).
@@
Assuming the slow-roll conditions are satisfied, we can immediately obtain the two slow-roll equations (i.e., Eqs. (2.14)) as
$H\simeq\frac{M}{2}\left(1-e^{-\sqrt{2/3}\phi}\right),\quad\phi_{N}\simeq{-2}\sqrt{\frac{2}{3}}\frac{e^{-\sqrt{2/3}\phi}}{\left(1-e^{-\sqrt{2/3}\phi}\right)},$ (3.12)
and the slow-roll parameters can be expressed in terms of the field $\phi$ as (i.e., Eqs. (2.15)):
@@
$\epsilon_{1}\simeq\frac{4e^{-2\sqrt{2/3}\phi}}{3\left(1-e^{-\sqrt{2/3}\phi}\right)^{2}},\qquad\quad\epsilon_{2}\simeq\frac{8e^{-\sqrt{2/3}\phi}}{3\left(1-e^{-\sqrt{2/3}\phi}\right)^{2}}.$ (3.13)
@@
With the help of Eq. (3.12), one can easily solve the scalar field solution in terms of the e-folding number $N_{k},$ and the solutions of the slow-roll parameters at the leading order can be obtained as:
$\epsilon_{1}\simeq\frac{3}{4N_{k}^{2}},\qquad\quad\epsilon_{2}\simeq\frac{2}{N_{k}},$ (3.14)
@@
and, using the leading order slow-roll approximation, i.e., Eqs. (2.16), we can quickly evaluate the scalar spectral index and the tensor-to-scalar ratio in terms of the e-folding number as:
$n_{s}\simeq 1-\frac{2}{N_{k}},\qquad\quad r\simeq\frac{12}{N_{k}^{2}}.$ (3.15)
@@
Figure 2: We plot the duration of reheating $N_{\rm re}$ and reheating temperature $T_{\rm re}$ given in Eqs. (3) and (3.9) as function of the scalar spectral index $n_{s}$ given by Eq. (3.15) parametrically for the case of Starobinsky inflation. Please note that different colors represent dynamics corresponding to different effective EoS parameter $w_{\rm re}$ as indicated in the figure. The blue-shaded region represents the $1\sigma$ constraint on the value of $n_{s}$ using ongoing observations [32, 33, 34, 35] with $n_{s}=0.9672\pm 0.0059$. The dark blue region shows the future projected bound on $n_{s}$ with a sensitivity of $10^{-3}$, assuming its central value remains unchanged. The temperature below the lighter red region is excluded due to the constraint from the electroweak scale, which is taken to be $100$ GeV.
@@
Figure 3: We plot the inflationary observables, tensor-to-scalar ratio ($r$) as a function of scalar spectral index ($n_{s}$) for the Starobinsky model using slow-roll approximations given by Eq. (3.15). This evolution is presented considering the bound on duration $N_{k}$ using the reheating regime. We consider different EoS parameters during reheating $w_{\rm re}$ and observe the constraint on the duration $N_{\rm k}$. Correspondingly, in the figure, blue line corresponds to the bound for $w_{\rm re}<1/3$, the gray line for $w_{\rm re}>1/3$ and the purple line for the future observational bound of $n_{\rm s}$. The blue-shaded region represents the $1\sigma$ constraint on the value of $n_{\rm s}$ using ongoing observations [32, 33, 34, 35] with $n_{\rm s}=0.9672\pm 0.0059$. The dark blue region shows the future projected bound on $n_{\rm s}$ with a sensitivity of $10^{-3}$, assuming its central value remains unchanged.
@@
Using the above equation along with Eqs. (3) and (3.9), one can obtain parametric dependence of $N_{\rm re}$ and $T_{\rm re}$ on $n_{s}$. In the case of the Starobinsky model, this is shown in Fig. 2. Keep in mind that, $N_{\rm re}\geq 0$, with $N_{\rm re}\ll 1$ is referred to as instantaneous reheating [92, 91]. It immediately leads to the constraint on the value of $n_{s}.$ For example, considering that during reheating, the evolution of the Universe effectively behaves as matter-expansion, i.e., $w_{\rm re}=0$. Thus, in the $1\sigma$ domain, we get the bound on $N_{k}$ as
$52\leq N_{k}\leq 56$ (3.16)
@@
$7\leq N_{\rm re}\leq 14,~{}~{}~{}~{}~{}\quad 10^{9}\,\text{GeV}\geq T_{\rm re}\geq 10^{6}\,\text{GeV}.$ (3.21)
@@
4 Numerical improvements
In the previous section, we discussed the CMB constraints on various inflationary models, which depend on the effective EoS parameter $w_{\rm re}$ as well as the duration of the reheating epoch $N_{\rm re}$. There are two issues associated with this analysis. First, around and at the end of inflation, the slow-roll approximation breaks down. Yet, in evaluating duration $N_{k}$ as well as energy density at end of inflation $\rho_{\rm end}$ in Eq. (3), we rely upon slow-roll approximations. Second, in evaluating perturbed observables in Eqs. (2.16), we consider only the leading-order contributions and ignore higher-order contributions, assuming their impact to be negligible. However, as anticipated in. Refs. [71, 72], considering accurate dynamics rather than the slow-roll approximations may lead to notable enhancements in these predictions. Additionally, one can consider the onset of reheating not at the end of inflation but at the bottom of the potential, as will be discussed in the subsequent section. The implementation and, therefore, the resulting corrections due to these considerations can be categorized into three parts:

### v3

@@ Abstract
The inflationary paradigm not only addresses early universe puzzles but also aligns well with the observational constraints, with slow-roll inflationary models fitting the best. Evaluating these model predictions requires considering both slow-roll inflationary dynamics and the subsequent reheating epoch. This involves the quantitative analysis that takes into account the effective equation of state (EoS) and duration of reheating, connecting these with the perturbations generated deep during the inflationary era. Given the complexities involved, many approximations are often used for simplification. However, as future observations are expected to improve the accuracy of these observables significantly, this work takes a different approach. Instead of relying on approximations, and instead of looking into the complex effects of (pre-)reheating, we focus on the corrections arising purely from more accurate analytical evaluations of the perturbations generated during the inflationary era itself. This is because the reheating dynamics is model-dependent and lack a single concrete analytical description, and thus introduce large uncertainties, making robust predictions difficult. In this article, we mainly incorporate two improvements: the first is the accurate dynamics of the slow-roll evolution, and, thus, the end of inflation; and the second is the higher-order slow-roll corrections to the perturbed observables. Our findings indicate that, by implementing these corrections, the theoretical predictions improve significantly. It also indicates that seemingly minor corrections can have significant effects on the perturbed observables and these refined predictions can be compared with future observations to potentially rule out models and help resolve the degeneracy problem of the inflationary paradigm.
The Early Universe, Inflationary Paradigm, Reheating, CMBR, Perturbations.
I Introduction
@@
Theoretical constraints derived from an inflationary model, however, not only depend on the dynamics of inflation itself but also depend on the reheating era, and therefore, studying reheating dynamics is crucial for accurately predicting the early universe behavior. There are two main approaches to this: the first is the qualitative analysis [42, 36, 37, 38, 39, 41, 40, 43, 46] which examines the exact decay process of the inflaton field through the parametric resonance, leading to the understanding of the microphysics involved during the reheating epoch. The second approach, which is the main focus of this manuscript, is the quantitative analysis [59, 54, 60, 44, 61, 45, 47, 49, 62, 52, 63, 64]. This approach looks at the macrophysics of the reheating epoch by constraining the duration ($N_{\rm re}$) and the equation of state (EoS) parameter ($w_{\rm re}$) during the reheating era using perturbations generated deep in the early (inflation) universe.
@@
Improving the analysis can again be performed in two main ways. The first one involves a detailed and careful consideration of the reheating analysis that can accurately depict the reheating parameters ($N_{\rm re},w_{\rm re}$). However, since it entails dealing with various distinct microphysical processes — requiring largely model-dependent treatment known as the qualitative analysis — and consequently introduces substantial uncertainties. For this reason, we have opted not to pursue this approach as it is beyond the scope of this work. Instead, we focus on the second, more straightforward improvement: the slow-roll methodology and approximations with a brief and simplest quantitative reheating epoch with constant $w_{\rm re}$. Recent findings in Refs. [72, 73] indicate that a near-accurate analytical solution can significantly enhance background predictions near the end of inflation, thereby improving $N_{k}$ [72]. For example, in the case of Starobinsky inflation, the improvement $\Delta N_{k}\sim 1.5$ suggests that
@@
The article is structured as follows: the following section II presents the model and the subsequent dynamics for the early universe dynamics. In this section, we also demonstrate the slow-roll conditions as well as the slow-roll approximations. In the subsequent section III, we study the quantitative analysis of reheating linking the inflationary perturbed variables to the observable constraints. This section, as an example, also examines the popular Starobinsky model of inflation, and with the help of the reheating analysis along with leading-order slow-roll approximations, we carefully obtain the observational consequences of the model. In Sec. IV, we highlight the dire need for the corrections to the dynamics, subsequently implement these changes, and obtain the improved constraint of the model, which turns out to be significant. Finally, we conclude our work in Sec. V.
A brief explanation about some terms and symbols we’ll be using. In our work, we use natural units where $\hbar=c=k_{B}=1$. We also define the reduced Planck mass as $M_{\rm pl}\equiv(8\pi G)^{-1/2}=1$. The metric signature we use is (,+,+,+)(-,+,+,+). Greek indices are contracted with the metric tensor $g_{\mu\nu}$, while Latin indices will be contracted with the Kronecker delta $\delta_{ij}$. We use symbols $\partial$ and $\nabla$ for partial and covariant derivatives, respectively. Additionally, the dots and primes above symbols represent derivatives for cosmic and conformal time, respectively, related to the Friedmann-Lemaître-Robertson-Walker (FLRW) line-element.
II General equations and slow-roll conditions
@@
$S=\frac{1}{2}\int{\rm d}^{4}x\sqrt{-g}(R-g^{\mu\nu}\partial_{\mu}\phi\partial_{\nu}\phi-2V(\phi)).$ (1)
@@
$T_{\mu\nu{(\phi)}}=\partial_{\mu}\phi\ \partial_{\nu}\phi-g_{\mu\nu}\left(\frac{1}{2}\partial_{\lambda}\phi\ \partial^{\lambda}\phi+V(\phi)\right).$ (4)
@@
${\rm d}s^{2}=-{\rm d}t^{2}+a^{2}(t){\rm d}\mathbf{x}^{2},$ (5)
where, $a(t)$ is the scalar factor, Eqs. (2) and (3) can be reduced to
$3H^{2}=\frac{1}{2}\dot{\phi}^{2}+V(\phi),$ (6)
$\dot{H}=-\frac{1}{2}\dot{\phi}^{2},$ (7)
$\ddot{\phi}+3H\dot{\phi}+V_{,\phi}=0.$ (8)
@@
$H=\sqrt{\frac{V}{3-\frac{1}{2}\phi_{N}^{2}}},\qquad\frac{H_{N}}{H}=-\frac{1}{2}\phi_{N}^{2},$ (9)
$\phi_{NN}+\left(3-\frac{1}{2}\phi^{2}_{N}\right)\left(\phi_{N}+\frac{V_{\phi}}{V}\right)=0,$ (10)
@@
$\mathcal{P}_{\mathcal{R}}=\frac{k^{3}}{2\pi^{2}}|\mathcal{R}_{k}|^{2}\equiv A_{\mathcal{R}}\left(\frac{k}{k_{*}}\right)^{n_{s}-1},\quad\mathcal{P}_{T}=2.\frac{k^{3}}{2\pi^{2}}|h_{k}|^{2}\equiv A_{T}\left(\frac{k}{k_{*}}\right)^{n_{T}}$ (11)
@@
$\epsilon_{\rm{1}}\equiv-\frac{\dot{H}}{H^{2}}=\frac{1}{2}\phi_{N}^{2},\qquad\epsilon_{\rm{2}}\equiv\frac{\dot{\epsilon_{\rm{1}}}}{H\epsilon_{\rm{1}}}=2\frac{\phi_{NN}}{\phi_{N}}.$ (12)
@@
$\epsilon_{\rm{1}}\ll 1,\quad\epsilon_{\rm{2}}\ll 1.$ (13)
These two condition are also referred to as slow-roll conditions. Under this conditions, Eqs. (9) and (10) become
$H^{2}\simeq\frac{1}{3}V,\quad\phi_{N}\simeq-\frac{V_{\phi}}{V},$ (14)
@@
$\epsilon_{1}\simeq\frac{1}{2}\left(\frac{V_{\phi}}{V}\right)^{2},\quad\epsilon_{2}\simeq 2\left(\frac{V^{2}_{\phi}}{V^{2}}-\frac{V_{\phi\phi}}{V}\right).$ (15)
Eqs. (14) are referred to as the slow-roll equations. By solving these simplified equations, one can obtain the approximate analytical solution of the Hubble parameter as well as the slow-roll parameters, and thus the dynamics of the universe in slow-roll regime.
@@
$A_{\mathcal{R}}\simeq\frac{H^{2}}{8\pi^{2}\epsilon_{1}},~~~~\quad n_{s}\simeq 1-2\epsilon_{1}-\epsilon_{2},~~~~\quad r\simeq 16\epsilon_{1}.$ (16)
@@
III Quantitative analysis of reheating
@@
This epoch involves intricate microphysical processes and is largely model-dependent. As a result, any analytical extension of its dynamics carries substantial uncertainties, making it extremely challenging to describe the epoch through a single, unified analytical framework. To circumvent this difficulty and achieve a model-independent treatment, a simplified quantitative approach has been developed in which the equation-of-state parameter during the entire reheating epoch, i.e., $w_{\rm re}$ is assumed to remain constant. Although this approximation does not capture the full microphysical complexity of reheating, it provides a useful macrophysical description that successfully connects inflationary predictions with observations and thereby enables broad, indirect constraints on the reheating scenario. In this case, the energy density during reheating effectively behaves as
$\rho\propto a^{-3(1+w_{\rm re})}.$ (17)
@@
$N_{\rm re}\equiv{\rm ln}\left(\frac{a_{\rm re}}{a_{\rm end}}\right)$ (18)
@@
$\frac{k}{a_{0}H_{0}}=\frac{a_{k}H_{k}}{a_{0}H_{0}}=\frac{a_{k}}{a_{\rm end}}\frac{a_{\rm end}}{a_{\rm{re}}}\frac{a_{\rm{re}}}{a_{\rm{eq}}}\frac{a_{\rm{eq}}}{a_{0}}\frac{H_{k}}{H_{\rm{eq}}}\frac{H_{\rm{eq}}}{H_{0}},$ (19)
@@
${\rm ln}\left(\frac{k}{a_{0}H_{0}}\right)=-N_{k}-N_{\rm re}-N_{\rm{RD}}+{\rm ln}\left(\frac{a_{\rm{eq}}H_{\rm{eq}}}{a_{0}H_{0}}\right)+{\rm ln}\left(\frac{H_{k}}{H_{\rm{eq}}}\right),$ (20)
@@
$g_{\rm{s,re}}\,T_{\rm{re}}^{3}=\left(\frac{a_{0}}{a_{\rm{re}}}\right)^{3}\left(2T_{0}^{3}+6\times\frac{7}{8}T_{\nu 0}^{3}\right),$ (22)
@@
$\frac{T_{\rm re}}{T_{0}}=\frac{a_{0}}{a_{\rm{eq}}}\frac{a_{\rm{eq}}}{a_{\rm{re}}}\left(\frac{43}{11g_{\rm{s,re}}}\right)^{1/3},$ (23)
and the Eq. (20) can be re-written as
@@
$T_{\rm{re}}=\frac{a_{0}T_{0}}{k}\left(\frac{43}{11g_{\rm{s,re}}}\right)^{1/3}H_{k}e^{-N_{\rm k}-N_{\rm{re}}}.$ (25)
@@
$H_{k}=\frac{\pi\sqrt{rA_{\mathcal{R}}}}{\sqrt{2}}.$ (26)
@@ III.1 Starobinsky inflation
$V(\phi)=\frac{3}{4}M^{2}\left(1-e^{-\sqrt{2/3}\phi}\right)^{2},$ (27)
@@
Figure 1: The Starobinsky potential given in Eq. (27).
@@
Assuming the slow-roll conditions are satisfied, we can immediately obtain the two slow-roll equations (i.e., Eqs. (14)) as
$H\simeq\frac{M}{2}\left(1-e^{-\sqrt{2/3}\phi}\right),\quad\phi_{N}\simeq{-2}\sqrt{\frac{2}{3}}\frac{e^{-\sqrt{2/3}\phi}}{\left(1-e^{-\sqrt{2/3}\phi}\right)},$ (28)
and the slow-roll parameters can be expressed in terms of the field $\phi$ as (i.e., Eqs. (15)):
@@
$\epsilon_{1}\simeq\frac{4e^{-2\sqrt{2/3}\phi}}{3\left(1-e^{-\sqrt{2/3}\phi}\right)^{2}},\qquad\quad\epsilon_{2}\simeq\frac{8e^{-\sqrt{2/3}\phi}}{3\left(1-e^{-\sqrt{2/3}\phi}\right)^{2}}.$ (29)
@@
With the help of Eq. (28), one can easily solve the scalar field solution in terms of the e-folding number $N_{k},$ and the solutions of the slow-roll parameters at the leading order can be obtained as:
$\epsilon_{1}\simeq\frac{3}{4N_{k}^{2}},\qquad\quad\epsilon_{2}\simeq\frac{2}{N_{k}},$ (30)
@@
and, using the leading order slow-roll approximation, i.e., Eqs. (16), we can quickly evaluate the scalar spectral index and the tensor-to-scalar ratio in terms of the e-folding number as:
$n_{s}\simeq 1-\frac{2}{N_{k}},\qquad\quad r\simeq\frac{12}{N_{k}^{2}}.$ (31)
@@
Figure 2: We plot the duration of reheating $N_{\rm re}$ and reheating temperature $T_{\rm re}$ given in Eqs. (III) and (25) as functions of the scalar spectral index $n_{s}$ given by Eq. (31) using leading order slow-roll approximations for the case of Starobinsky inflation. Please note that different colors represent dynamics corresponding to different effective EoS parameter $w_{\rm re}$ as indicated in the figure. The blue-shaded region represents the $1\sigma$ constraint on the value of $n_{s}$ using ongoing observations [32, 33, 34, 35] with $n_{s}=0.9672\pm 0.0059$. The dark blue region shows the future projected bound on $n_{s}$ with a sensitivity of $10^{-3}$, assuming its central value remains unchanged. The temperature below the lighter red region is excluded due to the constraint from the electroweak scale, which is taken to be 100100 GeV.
@@
Figure 3: We plot the inflationary observables, tensor-to-scalar ratio ($r$) as a function of scalar spectral index ($n_{s}$) for the Starobinsky model using leading order slow-roll approximations given by Eq. (31). This evolution is presented considering the bound on duration $N_{k}$ using the reheating regime. We consider different EoS parameters during reheating $w_{\rm re}$ and observe the constraint on the duration $N_{\rm k}$. Correspondingly, in the figure, blue line corresponds to the bound for $w_{\rm re}<1/3$, the gray line for $w_{\rm re}>1/3$ and the purple line for the future observational bound of $n_{\rm s}$. The blue-shaded region represents the $1\sigma$ constraint on the value of $n_{\rm s}$ using ongoing observations [32, 33, 34, 35] with $n_{\rm s}=0.9672\pm 0.0059$. The dark blue region shows the future projected bound on $n_{\rm s}$ with a sensitivity of $10^{-3}$, assuming its central value remains unchanged.
@@
Using the above equation along with Eqs. (III) and (25), one can obtain parametric dependence of $N_{\rm re}$ and $T_{\rm re}$ on $n_{s}$. In the case of the Starobinsky model, this is shown in Fig. 2. Keep in mind that, $N_{\rm re}\geq 0$, with $N_{\rm re}\ll 1$ is referred to as instantaneous reheating [93, 92]. It immediately leads to the constraint on the value of $n_{s}.$ For example, considering that during reheating, the evolution of the universe effectively behaves as matter-expansion, i.e., $w_{\rm re}=0$. Thus, in the $1\sigma$ domain, we get the bound on $N_{k}$ as
$52\leq N_{k}\leq 56$ (32)
@@
$7\leq N_{\rm re}\leq 14,~~~~~\quad 10^{9}\,\text{GeV}\geq T_{\rm re}\geq 10^{6}\,\text{GeV}.$ (37)
@@
IV Numerical improvements
In the previous section, we discussed the CMB constraints on various inflationary models, which depend on the effective EoS parameter $w_{\rm re}$ as well as the duration of the reheating epoch $N_{\rm re}$. There are two issues associated with this analysis. First, around and at the end of inflation, the slow-roll approximation breaks down. Yet, in evaluating duration $N_{k}$ as well as energy density at end of inflation $\rho_{\rm end}$ in Eq. (III), we rely upon slow-roll approximations. Second, in evaluating perturbed observables in Eqs. (16), we consider only the leading-order contributions and ignore higher-order contributions, assuming their impact to be negligible. However, as anticipated in. Refs. [72, 73], considering accurate dynamics rather than the slow-roll approximations may lead to notable enhancements in these predictions. Additionally, one can consider the onset of reheating not at the end of inflation but at the bottom of the potential, as will be discussed in the subsequent section. The implementation and, therefore, the resulting corrections due to these considerations can be categorized into three parts: